# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.2 Data Sets
Ranasinghe et al. [140] created a benchmark dataset they named SOLD [^67] https://huggingface.co/datasets/sinhala-nlp/SOLD: Sinhala Offensive Language Dataset which contains 10,000 posts from Twitter annotated both at sentence-level and token-level with the two classes offensive and not offensive. In the same paper, they also introduce the dataset SemiSOLD [^68] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach. The relevant code is also available on Github 696969%****␣03_Sinhala.tex␣Line␣75␣****https://github.com/Sinhala-NLP/SOLD. The subsequent work by Dmonte et al. [141] used the SOLD data set in the evaluation process of their new MT-Offense [^70] https://github.com/LanguageTechnologyLab/MT-Offense data set which they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
@@
Singh et al. [159] introduced the Aya Data set [^87] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^88] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^89] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality. Kudugunta et al. [160] introduced MADLAD-400909090https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents. They have also released the trained checkpoints on GitHub [^91] https://github.com/google-research/google-research/tree/master/madlad_400. A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al. [161] along with data [^92] https://huggingface.co/datasets/afaji/cvqa and a leader board [^93] https://eval.ai/web/challenges/challenge-page/2305/overview. Pushpananda et al. [162] claims to have created a manually annotated parallel corpus named TamSiPara with 25k parallel sentences in the Si-Ta direction and 22k parallel sentences in the Ta-Si direction. But it is not freely and openly accessible. While the stated purpose of Ranasinghe [163] is creating a tokenizer for Sinhala and not creating a text classification data set, the repository does contain [^94] https://github.com/Buddhilive/sltk/blob/main/data/sin.csv a data set of 26k Sinhala news titles with classification labels. A visual Q&A benchmark data set named ALM-Bench [^95] https://huggingface.co/datasets/MBZUAI/ALM-Bench was created by Vayani et al. [164] to evaluate Q&A performance of LLMs on 100 languages including Sinhala. They initially machine translated LLaVA-Bench (In-the-Wild) dataset [165] data set and then extended it by searching for images under country name, language name, and cultural category on the web so that they can include culturally relevant questions and answer for each language. In their own experiment results, they report that GPT-4o [166] performance on Sinhala is superior to other tested LLMs on this data set. Ranathunga et al. [167] has created a text simplification dataset named SiTSE [^96] https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation. It has 1000 complex Sinhala texts along with two separate simplifications for each of them. Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data. Ranathunga et al. [168] created multiNER [^97] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil. It contains 99k tagged Sinhala words. They have used PER, ORG, LOC, and MISC tags as well as O. Singh et al. [169] introduced Global-MMLU [^98] https://huggingface.co/datasets/CohereForAI/Global-MMLU, a multilingual extension of the original English Massive Multitask Language Understanding (MMLU) dataset [170]. They claim to have improved the data set beyond what Üstün et al. [171] did before. Chavinda and Thayasivam [172] introduces two publicly available Sinhala hate speech data sets. One is from Facebook [^99] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter100100100https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech. A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released101101101https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [173]. One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish. Burchell et al. [174] introduced HPLT v2102102102https://huggingface.co/datasets/HPLT/HPLT2.0_cleaned a multilingual monolingual and parallel corpora that includes 273k Sinhala Sentences. These sentences were extracted from the Internet Archive(IA) and the Common Crawl (CC). They have also publicly released their translation models103103103https://huggingface.co/collections/HPLT/hplt-20-uni-direction-translation-models-67f2fc7ae54845f9b182957a.
@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [175]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [176, 177, 178, 179, 180, 181] as well. The dictionary by Kulatunga [182] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [183] which bears a more than passing resemblance to the above dictionary by Kulatunga [182]Hettige and Karunananda [79] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [184]. The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github104104104https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara105105105http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [187] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [37] to information extraction based on semantic similarity [188, 189] or semantic oppositeness [190]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [48]. For a time it was hosted on [191] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [192] have cloned to use in their application uploaded on github106106106https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [193] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [48].
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [206], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [208] which then extended to a Hidden Markov Model (HMM) based approach [209] and an analysis of unknown words [210, 211]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [212]. A RESTFul PoS tagging web service created by Jayaweera and Dias [213] using the above research can still be accessed107107107http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [214]. The study by Kothalawala et al. [215] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [216] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [217] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [115]. The study by Sathsarani et al. [218] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [118] is available to download at GitHub 108108108http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [223] and by Senevirathne et al. [224] are not accessible to the public. Azeez and Ranathunga [225] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [118] and PoS tagging [117]Anuruddha [226] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [227], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [228] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [229] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [230] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). A data set for Sinhala NER named multiNER109109109https://github.com/suralk/multiNER was introduced by Ranathunga et al. [168]. It includes parallel NER data for Sinhala, Tamil, and English. Rathnayake and Asanka [231] simply uses LLaMA for Sinhala NER.
@@ 3.9 Semantic Similarity
A Sinhala semantic similarity measure has been developed for short sentences by Kadupitiya et al. [232]. This work has been then extended by Kadupitiya et al. [233] for the application use case of short answer grading. Data and tools for these projects are not publicly available. A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [234] where they demonstrate their results for Sinhala and Tamil. A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al. [235]. Deepal et al. [236] proposed to use Siamese network which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity. They report an F1 of 0.9041 for their model calculated by cosine distance. They discuss how they used a number of publicly available data sets and refined them to be used in their task, nevertheless, they do not publicly release the resulting data set.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [237]. A smaller implementation of Sinhala news classification has been attempted by de Silva [28]. As mentioned in Section 3.2, their news corpus is publicly available110110110https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [238]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [239] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [240] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [241] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [242, 243] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [244]Demotte and Ranathunga [245] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [246]. The attempt by Sameemdeen and Selvanthan [247] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [241]Gunasekara and Haddela [242]Lakmali and Haddela [238], and Buddhika et al. [248]. Then they propose active learning [249, 250] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [251] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [37]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [252]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [253]. The study by [254] used Genetic Algorithm on Lucene111111111https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [255] uses adapter-based [256, 257, 258, 259, 260, 261] fine-tuning on XLM-R [262], for code-mixed and code-switched text. Kirindage and Godewithana [263] used LDA [264] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [265] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them into the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [107]. Both their code112112112https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1131131133https://tinyurl.com/42un7a9y, Sinhala News source classification data set114114114https://bit.ly/3PBTyfo, Sinhala News category classification data set115115115https://bit.ly/3AsfK7u, and Sinhala writing style classification data set116116116https://bit.ly/3PARuUS) are publicly available. Faumi et al. [136] also worked on Sinhala writing style identification (author identification) using their own data set117117117https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text. The study by Wijayarathna and Jayalal [266] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [267] used capsule-based methods recommended by Senevirathne et al. [246] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [268] compared Word2Vec, FastText and Doc2vec [269] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [270] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter. The work by Gunawardana and Thamarasee [271] attempted to predict the subsequent popularity of Sinhala language YouTube videos by analysing a number of associated factors. Ilangeshwaran et al. [272] introduced Saliency-based token Swap (SSwap), a data augmentation technique to improve Sinhala text classification. The work by Rathnayake and Asanka [231] uses LLaMA for Named entity classification in Sinhala documents and then runs Fuzzy C-Means clustering on it. Next, they utilise LLaMA again to propose names to the created clusters.

### v26

@@ 3.2 Data Sets
Ranasinghe et al. [143] created a benchmark dataset they named SOLD [^67] https://huggingface.co/datasets/sinhala-nlp/SOLD: Sinhala Offensive Language Dataset which contains 10,000 posts from Twitter annotated both at sentence-level and token-level with the two classes offensive and not offensive. In the same paper, they also introduce the dataset SemiSOLD [^68] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach. The relevant code is also available on Github  [^69] https://github.com/Sinhala-NLP/SOLD. The subsequent work by Dmonte et al. [144] used the SOLD data set in the evaluation process of their new MT-Offense [^70] https://github.com/LanguageTechnologyLab/MT-Offense data set which they created by using NMT to translate the English offensive language benchmark dataset, OLID [145].
@@
Singh et al. [162] introduced the Aya Data set [^87] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^88] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^89] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality. Kudugunta et al. [163] introduced MADLAD-400909090https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents. They have also released the trained checkpoints on GitHub [^91] https://github.com/google-research/google-research/tree/master/madlad_400. A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al. [164] along with data [^92] https://huggingface.co/datasets/afaji/cvqa and a leader board [^93] https://eval.ai/web/challenges/challenge-page/2305/overview. Pushpananda et al. [165] claims to have created a manually annotated parallel corpus named TamSiPara with 25k parallel sentences in the Si-Ta direction and 22k parallel sentences in the Ta-Si direction. But it is not freely and openly accessible. While the stated purpose of Ranasinghe [166] is creating a tokenizer for Sinhala and not creating a text classification data set, the repository does contain [^94] https://github.com/Buddhilive/sltk/blob/main/data/sin.csv a data set of 26k Sinhala news titles with classification labels. A visual Q&A benchmark data set named ALM-Bench [^95] https://huggingface.co/datasets/MBZUAI/ALM-Bench was created by Vayani et al. [167] to evaluate Q&A performance of LLMs on 100 languages including Sinhala. They initially machine translated LLaVA-Bench (In-the-Wild) dataset [168] data set and then extended it by searching for images under country name, language name, and cultural category on the web so that they can include culturally relevant questions and answer for each language. In their own experiment results, they report that GPT-4o [169] performance on Sinhala is superior to other tested LLMs on this data set. Ranathunga et al. [170] created multiNER [^96] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil. It contains 99k tagged Sinhala words. They have used PER, ORG, LOC, and MISC tags as well as O. Singh et al. [171] introduced Global-MMLU [^97] https://huggingface.co/datasets/CohereForAI/Global-MMLU, a multilingual extension of the original English Massive Multitask Language Understanding (MMLU) dataset [172]. They claim to have improved the data set beyond what Üstün et al. [173] did before. Chavinda and Thayasivam [174] introduces two publicly available Sinhala hate speech data sets. One is from Facebook [^98] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter [^99] https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech. A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released100100100https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [175]. One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish. By further manually tagging this dataset, De Mel and de Silva [176] introduced GeeSanBhava101101101https://bit.ly/SinhalaYoutubeComments, the first Sinhala sentiment analysis dataset tagged using Russell’s Valence-Arousal model [177] as opposed to arbitrary sentiment classes. Burchell et al. [178] introduced HPLT v2102102102https://huggingface.co/datasets/HPLT/HPLT2.0_cleaned a multilingual monolingual and parallel corpora that includes 273k Sinhala Sentences. These sentences were extracted from the Internet Archive(IA) and the Common Crawl (CC). They have also publicly released their translation models103103103https://huggingface.co/collections/HPLT/hplt-20-uni-direction-translation-models-67f2fc7ae54845f9b182957a. Chang et al. [179] created Global PIQA104104104https://mrlbenchmarks.github.io/, a multilingual extension to PIQA [180] in which they include physical commonsense reasoning triples covering 100+ languages such that the examples are culturally aligned to the relevant languages. The objective of this dataset is to function as a benchmark to measure the cultural alignment of LLMs. Senaratna [181] has created and perpetually maintains Sri Lanka Document Datasets in which he curates data sets which include Sinhala text such as: Hansards105105105https://github.com/nuuuwan/lk_hansard/, Police Press Releases106106106https://github.com/nuuuwan/lk_police_press_releases/, Legal Documents107107107https://github.com/nuuuwan/lk_legal_docs/, Cabinet Decisions108108108https://github.com/nuuuwan/lk_cabinet_decisions/, Treasury Press Releases109109109https://github.com/nuuuwan/lk_treasury/, Sri Lanka Presidential Media Division press releases110110110https://github.com/nuuuwan/lk_pmd/, News111111111https://github.com/nuuuwan/lk_news/, and Educational Publications from the Educational Publications Department112112112https://github.com/nuuuwan/lk_edupub/. A collection of Sinhala-Sinhala, Sinhala-English, English-Sinhala, and Pali-Sinhala dictionaries is available on GitHub113113113https://github.com/pnfo/arutha.lk/tree/main/dev/dict-input courtesy of aruth.lk. Another Pali-Sinhala dictionary is available on Github114114114https://github.com/pnfo/pali-sinhala-dictionary by the Path Nirvana project. The same project provides a large repository of Buddhist books with Sinhala and Pali text 115115115https://github.com/pathnirvana/pitaka.lk/tree/master/books116116116https://github.com/pathnirvana/sinhala-tools/tree/master/dev/buddhist-books/text. Further, they have some News data collected from Silumina117117117https://github.com/pathnirvana/sinhala-tools/tree/master/dev/silumina, Mawbima118118118https://github.com/pathnirvana/sinhala-tools/tree/master/dev/mawbima, and AdaDerana119119119https://github.com/pathnirvana/sinhala-tools/tree/master/dev/adaderana. An associated profile contains a large Sinhala TTS data set120120120https://github.com/pnfo/sinhala-tts-dataset consisting of 6248 sentences with 13.8 hours of recordings over two speakers: one male and one female. A wiki version of the Sinhala Encyclopedia maintained by the Sri Lankan government is available online121121121https://t.co/k8gFT9puM0. It seems it has not been updated recently. However, some of its articles are of higher quality than the same of Sinhala Wikipedia122122122https://si.wikipedia.org/. Google has made available a comprehensive IPA conversion (pronunciation) dictionary of a large Sinhala vocabulary123123123https://raw.githubusercontent.com/google/language-resources/master/si/data/lexicon.tsv. A PoS tagged data set of 250,000+ Sinhala words is available through Google Groups124124124https://t.co/jpQz74YEdQ. A rudimentary Sinhala Tokenizer, Stemmer, and a PoSTagger have been published on GitHub125125125https://github.com/ysenarath/sinling by the user ysenarath. While the research paper by Ekanayaka et al. [182] seems to be lost, the news corpus created for that study, which contains 1000 labelled and 50,000 unlabelled Sinhala news reports collected from various sources, has been published to github126126126https://github.com/rksk/sinhala-news-analysis/tree/master/sinhala-news-corpus. The repository also contains some code developed for the relevant work 127127127https://github.com/rksk/sinhala-news-analysis/.
@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [183]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [184, 185, 186, 187, 188, 189] as well. The dictionary by Kulatunga [190] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [191] which bears a more than passing resemblance to the above dictionary by Kulatunga [190]Hettige and Karunananda [82] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [192]. The study by Wickramasinghe and De Silva [150] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Si pairs. Both their code and the two versions of the dictionary are publicly available on github128128128https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara129129129http://crawler.nlpc.uom.lk/ was built by Iniyage et al. [193]. It is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [196] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [39] to information extraction based on semantic similarity [197, 198] or semantic oppositeness [199]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [50]. For a time it was hosted on [200] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [201] have cloned to use in their application uploaded on github130130130https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [202] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [50].
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [215], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [217] which then extended to a Hidden Markov Model (HMM) based approach [218] and an analysis of unknown words [219, 220]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [221]. A RESTFul PoS tagging web service created by Jayaweera and Dias [222] using the above research can still be accessed131131131http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [223]. The study by Kothalawala et al. [224] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [225] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [226] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [118]. The study by Sathsarani et al. [227] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.
@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [121] is available to download at GitHub 132132132http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [232] and by Senevirathne et al. [233] are not accessible to the public. Azeez and Ranathunga [234] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [121] and PoS tagging [120]Anuruddha [235] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [236], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [237] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [238] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [239] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). A data set for Sinhala NER named multiNER133133133https://github.com/suralk/multiNER was introduced by Ranathunga et al. [170]. It includes parallel NER data for Sinhala, Tamil, and English. Rathnayake and Asanka [240] simply uses LLaMA for Sinhala NER. Sinhala NER is discussed along with other Indic languages in the survey by Deshmukh et al. [241].
@@ 3.9 Semantic Similarity
A Sinhala semantic similarity measure has been developed for short sentences by Kadupitiya et al. [242]. This work has been then extended by Kadupitiya et al. [243] for the application use case of short answer grading. Data and tools for these projects are not publicly available. A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [244], where they demonstrate their results for Sinhala and Tamil. A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al. [245]. Deepal et al. [246] proposed to use a Siamese network, which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity. They report an F1 of 0.9041 for their model, calculated by cosine distance. They discuss how they used a number of publicly available data sets and refined them to be used in their task, nevertheless, they do not publicly release the resulting data set. Ranasinghe et al. [247] introduced MUSTS134134134https://github.com/TharinduDR/MUSTS, a multilingual semantic similarity benchmark that includes Sinhala. They use the Sinhala short sentence similarity data set created by Kadupitiya et al. [243] as the Sinhala source for this.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [248]. A smaller implementation of Sinhala news classification has been attempted by de Silva [30]. As mentioned in Section 3.2, their news corpus is publicly available135135135https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [249]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [250] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [251] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [252] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [253, 254] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [255]Demotte and Ranathunga [256] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [257]. The attempt by Sameemdeen and Selvanthan [258] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [252]Gunasekara and Haddela [253]Lakmali and Haddela [249], and Buddhika et al. [259]. Then they propose active learning [260, 261] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [262] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [39]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [263]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [264]. The study by [265] used Genetic Algorithm on Lucene136136136https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [266] uses adapter-based [267, 268, 269, 270, 271, 272] fine-tuning on XLM-R [273], for code-mixed and code-switched text. Kirindage and Godewithana [274] used LDA [275] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [276] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them into the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [110]. Both their code137137137https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1381381383https://tinyurl.com/42un7a9y, Sinhala News source classification data set139139139https://bit.ly/3PBTyfo, Sinhala News category classification data set140140140https://bit.ly/3AsfK7u, and Sinhala writing style classification data set141141141https://bit.ly/3PARuUS) are publicly available. Faumi et al. [139] also worked on Sinhala writing style identification (author identification) using their own data set142142142https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text. Ranathunga et al. [277] also worked on writing style identification; however their work is specifically on classifying Sinhala answers as AI-generated or human-written. They reported 86% accuracy for LSTM running on TF-IDF features.
The study by Wijayarathna and Jayalal [278] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [279] used capsule-based methods recommended by Senevirathne et al. [257] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [280] compared Word2Vec, FastText and Doc2vec [281] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [282] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter. The work by Gunawardana and Thamarasee [283] attempted to predict the subsequent popularity of Sinhala language YouTube videos by analysing a number of associated factors. Ilangeshwaran et al. [284] introduced Saliency-based token Swap (SSwap), a data augmentation technique to improve Sinhala text classification. Wickramasinghe et al. [285] proposed a Sinhala news personalisation model built on SinBERT for classification. The work by Rathnayake and Asanka [240] uses LLaMA for Named entity classification in Sinhala documents and then runs Fuzzy C-Means clustering on it. Next, they utilise LLaMA again to propose names to the created clusters.